OpenStudy (anonymous):

FACTORING 10x^2-39x-27

OpenStudy (zzr0ck3r):

\[ax^2+bx+c\\10*(-27) =-270=a*c\\ 45*(-6)=-210=a*c\\ \space 45+(-6) = -39=b \\so\\ 10x^2-39x-27=\\10x^2+6x-45x-27=\\2x(5x+3)-9(5x+3) =\\ (2x-9)(5x+3)\]

OpenStudy (anonymous):

THANK YOU!!